A Bibliography of Publications on Software Standards

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

12 December 2019 Version 1.17

Abstract

This bibliography records publications on national and international standards for computer software.

Title word cross-reference

101-1983 [Gra83]. 198x [Ame87a]. 199x [Ame90].

24-dots [Jap83]. 2nd [IEE88].

6234-1983 [Jap83]. 6626-1978 [Jap78].

754 [Hou81]. 754-1985 [P7585]. 770X3.97 [Ame83b]. 770X3.97-1983 [Ame83b].

= [Int88].

ACM [SIG77, SIG79]. ACM/SIGGRAPH [SIG77, SIG79]. Ada [Ame83a]. American [Ame89, Org91, Ame86a, Ame88a, Ame88b, Ame90, Sch90]. Annotated [Sch90]. ANS [X3J76]. ANSI [Ame89, Ame66, Ame78, Ame87a, Ame85, Ame86b, Ame83b, P7585, Sch90]. ANSI/IEEE [Ame83b, P7585]. API [POS90]. Application [POS90]. Applications [Hou81]. Arithmetic [CCG^{+}84, Coo80, Coo81b, Hou81, P7585, Ste81]. ASA [Hei66]. Author [Bry88]. authors [Smi87b].

been [Sum93]. Binary [P7585, Ste81].

C [Ame89, Jap78, Jap83, Ame86a, Ame88a, Ame88b, POS90, Pla92, Plu87, Sch90]. CD [Int88, IE88]. CD-ROM [Int88, IE88]. Character [Jap78, Jap83, The91]. Clarification [X3J69, X3J71]. Code [Jap78]. Color [Cor89]. Committee [SIG77, SIG79]. compacts [Int88]. Computer [Ame85, Ame86b, Ame88c, EKP84, EKP87, IEE86, IEE88]. Conference [IEE88].

Denormalized [Coo81a]. Description [Ame86b, Ame87b, Smi92]. destinés [Int88]. Development [Hei66]. d’information [Int88]. disques [Int88]. Document [Gra83, Smi92]. Dot [Jap83]. dots [Jap83]. Draft [Ame86a, Ame88a, Ame88b, X3J76, Ame87a, Ame90, Ame88c, Plu87].

échange [Int88]. Electronic [Org91]. Encoding [The91, Cor89]. Environments [IEE86]. Errata [Coo81b]. Extended [Ame90, Sum93].

fichier [Int88]. file [Int88, IE88]. fixe [Int88]. Floating [CCG^{+}84, Coo80, Coo81b, Hou81, P7585, Ste81]. Floating-Point [CCG^{+}84, P7585, Ste81]. Fortran [Ame66, X3J69, X3J71, X3J76, Ame78, Ame87a, Ame90, Hei66]. Functional [Ame87b].

GCA [Gra83]. GenCode [Gra83]. Generalized [Bry88, Fee88, Gra83, Smi87a, Smi87b]. GKS [Ame85, EKP84, EKP87]. Graphic [SIG77, SIG79]. Graphical [Ame85, Ame88c]. Graphics [Ame85, Ame86b, Ame88c, EKP84, EKP87, Jap78]. Guide [Bry88, Coo80, Coo81b, SS88, Zlo91, Lew91]. guidelines [Smi87b].

Hierarchical [Ame88c]. History [Hei66].

IEEE [Ame83b, Hou81, IEE85, IEE86, P7585, IEE88]. IEEE-754 [Hou81]. Implementation [Coo80, Coo81b]. implemented [Sum93]. independent [CCG^{+}84]. Industrial [Jap78, Jap83]. Information [Ame85, Ame86b, Ame88c, Int88, IE88, Jap78]. Initial [X3J69]. Interactive [Ame88c]. interchange [Int88, IE88, Jap78]. Interface [POS90, POS93a, POS93b]. Interfaces [IEE85]. Introduction [Sum93]. ISO [JW85, SS88, Sum93]. Issues [Smi87a].

Japanese [Jap78, Jap78, Jap83]. JIS

[Jap78, Jap83].

Kernel [Ame85].

Language [Ame83a, Ame86a, Ame88a, Ame88b, Ame89, Ame90, Ame83b, Fee88, Gra83, POS90, Smi87a, Smi87b, Bry88]. Languages [Sch90, Smi92]. length [CCG^{+}84]. Library [Pla92].

Manual [JW85]. Manuscript [Org91]. Markup [Bry88, Fee88, Gra83, Org91, Smi87a, Smi87b]. Matrix [Jap83]. mémoire [Int88]. Metafile [Ame86b]. Metalanguage [Gra83]. Microprocessor [IEE85]. Military [Ame83a].

National [Ame86a, Ame88a, Ame88b, Ame89, Ame90, Sch90]. Notes [Plu87]. Numbers [Coo81a].

Operating

[IEE85, IEE86, POS93a, POS93b]. optiques [Int88].

Part [POS93a, POS93b]. Pascal [Ame83b, JW85]. Patterns [Jap83]. PHIGS [Ame87b]. Picture [Ame86b]. Planning [SIG77, SIG79]. Point [CCG^{+}84, Coo80, Coo81b, Hou81, P7585, Ste81]. Portable [IEE86, POS93a, POS93b, Lew91]. POSIX [POS93a, POS93b, Lew91]. POSIX.1 [Lew91, Zlo91]. POSIX.4 [Gal95]. Preparation [Org91]. Printers [Jap83]. Proceedings [IEE88]. Processing [Smi92, Int88, IE88]. Program [POS90]. Programmer [Ame88c, Zlo91, Lew91]. Programming [Ame83a, Ame86a, Ame88a, Ame88b, Ame89, Ame90, Ame83b, EKP84, EKP87, Gal95, Sch90]. programs [Lew91]. Progress [X3J69]. proposal [Ame88c]. Proposed [Ame86a, Ame88a, Ame88b, X3J76, Ame87a, Ame90, CCG^{+}84, Coo80, Coo81b, Hou81, Ste81].

Radix [CCG^{+}84]. Radix- [CCG^{+}84]. Real [Gal95]. Related [Smi87a, Smi92]. Report [X3J71, JW85, SIG77, SIG79]. Revision [Ame87b]. ROM [Int88, IE88].

Second [X3J71]. Set [Jap78]. SGML [Bry88, Fee88, Gra83, SS88, Smi87b, Smi92]. Shell [POS93a, POS93b]. SIGGRAPH [SIG77, SIG79]. Specifications [IEE85]. Standard [Ame83a, Ame89, CCG^{+}84, Coo80, Coo81b, EKP84, EKP87, Fee88, Gra83, Hou81, IEE85, IEE86, P7585, Jap78, Jap83, JW85, Org91, Plu87, Sch90, Smi87a, Smi87b, Ste81, The91, Zlo91, Lew91, Ame86a, Ame88a, Ame88b, Ame90, Bry88, Pla92, Cor89]. Standardization [Hei66]. Standards [X3J69, X3J71, SIG77, SIG79, Smi92]. Status [SIG77, SIG79]. Storage [Ame86b]. structure [Int88, IE88]. Summary [Hei66]. System [Ame85, Ame88c, IEE86, POS90, POS93a, POS93b]. Systems [Ame85, Ame86b, Ame88c, IEE85].

Traitement [Int88]. Transfer [Ame86b]. Trial [IEE85, IEE86]. Trial-Use [IEE85, IEE86].

Underflow [Coo81a]. UNIX [Lew91]. Use [IEE85, IEE86]. User [JW85, SS88]. Utilities [POS93a, POS93b].

Version [The91]. Volume [POS93a, POS93b, Int88, IE88]. Volumes [The91].

Word [CCG^{+}84].

Word-length-independent [CCG^{+}84]. Workstations [IEE88]. World [Gal95]. Worldwide [The91]. writing [Lew91].

X3.122 [Ame86b]. X3.122-1986 [Ame86b]. X3.124 [Ame85]. X3.124-1985 [Ame85]. X3.144.1988 [Ame88c]. X3.159 [Ame89]. X3.159-1989 [Ame89]. X3.198 [Ame90]. X3.9 [Ame78, Ame87a, Ame66]. X3.9-1966 [Ame66]. Xerox [Cor89].

References

ANSI:ftn66

[Ame66] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9-1966, 1966. Approved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [X3J69] and [X3J71], and history [Hei66].

ANSI:ftn77

[Ame78] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9–1978, 1978. Approved April 3, 1978 (also known as Fortran 77). See also draft [X3J76].

ANSI:ada

[Ame83a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Military Standard Ada Programming Language, February 17 1983. Also MIL-STD-1815A.

ANSI:pascal

[Ame83b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. The Pascal Programming Language. ANSI/IEEE 770X3.97-1983, 1983.

ANSI:gks

[Ame85] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems—Computer Graphics— Graphical Kernel System (GKS). ANSI X3.124-1985, 1985. Includes Fortran bindings to GKS.

ANSI:c86

[Ame86a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, October 1 1986.

ANSI:metafile

[Ame86b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems—Computer Graphics—Metafile for the Storage and Transfer of Picture Description Information. ANSI X3.122-1986, 1986.

ANSI:ftn8x

[Ame87a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed ANSI Fortran X3.9–198x, September 18 1987.

ANSI:phigs+

[Ame87b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. PHIGS+ Functional Description, Revision 2.0, July 20 1987.

ANSI:c88

[Ame88a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, January 11 1988.

ANSI:c88b

[Ame88b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, October 31 1988.

ANSI:phigs

[Ame88c] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems—Computer Graphics—Programmer’s Hierarchical Interactive Graphical System. Draft proposal X3.144.1988, 1988.

ANSI:c89

[Ame89] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. American National Standard Programming Language C, ANSI X3.159-1989, December 14 1989.

ANSI:ftn9x

[Ame90] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language Fortran Extended X3.198–199x, September 24 1990.

Bryan:1988:SAG

[Bry88] Martin Bryan. SGML: An Author’s Guide to the Standard Generalized Markup Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17535-5. xvii + 364 pp. LCCN QA76.73.S44 B79 1988. UK£16.95.

Cody:IEEE-P854

[CCG^{+}84] William J. Cody, Jr., Jerome T.

Coonen, David M. Gay, K. Hanson, David Hough, W. Kahan, R. Karpinski, John F. Palmer, F. N. Ris, and D. Stevenson. A proposed radix- and word-length-independent standard for floatingpoint arithmetic. IEEE Micro, 4 (4):86–100, July/August 1984.

Coonen:fps-guide

[Coo80] Jerome T. Coonen. An implementation guide to a proposed standard for floating point arithmetic. Computer, 13(1):68–79, January 1980. See errata in [Coo81b].

Coonen:ufl-denorm

[Coo81a] Jerome T. Coonan. Underflow and the denormalized numbers. Computer, 14(3):75–87, March 1981.

Coonen:fps-guide-errata

[Coo81b] Jerome T. Coonen. Errata: An implementation guide to a proposed standard for floating point arithmetic. Computer, 14(3):62, March 1981. See also [Coo80].

Xerox:color

[Cor89] Xerox Corporation. The Xerox Color Encoding Standard. Technical Report XNSS 288811, Xerox Systems Institute, March 1989.

Enderle:CGP84

[EKP84] G. Enderle, K. Kansy, and G. Pfaff. Computer Graphics Programming. GKS—The Graphics Standard. Symbolic Computation, Editor: J. Encarnação and P. Hayes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-16317-4 (New York), 3-540-16317-4 (Berlin), 0-387-11525-0 (New York), 3-540-11525-0 (Berlin). xvi + 542 pp. LCCN T385 .E531 1984.

Enderle:CGP87

[EKP87] G. Enderle, K. Kansy, and G. Pfaff. Computer Graphics Programming. GKS—The Graphics Standard. Symbolic Computation, Editor: J. Encarnação and P. Hayes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1987. ISBN 0-387-11525-0, 3-540-11525-0. xxiii + 651 pp. LCCN T385 .E531 1987.

Feeney:SGM88

[Fee88] Mary Feeney. The Standard Generalized Markup Language (SGML). British Library Research and Development Dept. and Library & Information Technology Centre, London, UK, 1988. 15 pp. LCCN Z286.E43 F44 1988.

Gallmeister:1995:PPR

[Gal95] Bill Gallmeister. POSIX.4: Programming for the Real World. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1995. ISBN 1-56592-074-0. xviii + 548 pp. LCCN QA76.76.O63 G34 1995. US$29.95.

GCA:DMM83

[Gra83] Graphic Communications Association. GCA Standard 101-1983, Document Markup Metalanguage: GenCode and the Standard Generalized Markup Language (SGML). Graphic Communications Association, 100 Daingerfield Road, Alexandria, VA 22314-2804, USA, 1983. ISBN 0-89740-244-8. xxiv + 264 pp. LCCN QA76.9.T48 G33 1983. US$48.00. Adopted by Department of Defense, 10 August 1983.

Heising:ftn

[Hei66] W. P. Heising. History and summary of Fortran standardization development for the ASA. Communications of the Association for Computing Machinery, 7:590–625, 1966. See also final standard [Ame66].

Hough:fps-applications

[Hou81] David Hough. Applications of the proposed IEEE-754 standard for floating point arithmetic. Computer, 14(3):70–74, March 1981.

ISO:1988:IPVb

[IE88] International Organization for Standardization and European Computer Manufacturers Association. Information processing: volume and file structure of CD-ROM for information interchange. International standard; ISO 9660 ISO 9660: 1988 (E), International Organization for Standardization, Geneva, Switzerland, April 15, 1988. x + 31 pp.

IEEE:ITU85

[IEE85] IEEE. IEEE Trial-Use Standard Specifications for Microprocessor Operating Systems Interfaces. Wiley-Interscience, New York, NY, USA, 1985. ISBN 0-471-01073-1. 173 pp. LCCN QA76.5 .I23 1985. IEEE Std. 855.

IEEE:ITU86

[IEE86] IEEE. IEEE Trial-Use Standard Portable Operating System for Computer Environments. Wiley-Interscience, New York, NY, USA, IEEE Std 1003.1 edition, April 1986. ISBN 0-471-85027-6. 207 pp. LCCN TK 275 I5 Std 1003.1.

IEEE:workstations

[IEE88] Proceedings of the 2nd IEEE Conference on Computer Workstations. IEEE Computer Society, 345 E. 47th St, New York, NY 10017, USA, March 7–10 1988.

ISO:1988:IPVa

[Int88] International Organization for Standardization. Information processing: volume and file structure of CD-ROM for information interchange = Traitement de l’information: structure de volume et de fichier des disques optiques compacts a mémoire fixe (CD-ROM) destinés a l’échange d’information. International standard; ISO 9660. International Organization for Standardization, Geneva, Switzerland, corrected and reprinted. edition, 1988. x + 31 pp.

JISCII:charset

[Jap78] Japanese Standards Association, 1-24, Akasaka 4 Chome, Minatoku, Tokyo, 107 Japan. Japanese Industrial Standard JIS C 6626-1978 Code of the Japanese Graphics Character Set for Information Interchange, 1978.

JISCII:dot-printers

[Jap83] Japanese Standards Association, 1-24, Akasaka 4 Chome, Minatoku, Tokyo, 107 Japan. Japanese Industrial Standard JIS C 6234-1983 24-dots Matrix Character Patterns for Dot Printers, 1983.

Jensen:PUM85

[JW85] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report—ISO Pascal Standard. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1985. ISBN 0-387-96048-1, 3-540-96048-1. xvi + 266 pp. LCCN QA76.73.P2 J46 1985. Revised by Andrew B. Mickel and James F. Miner.

Lewine:1991:PPG

[Lew91] Donald A. Lewine. POSIX programmer’s guide: writing portable UNIX programs with the POSIX.1 standard. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-73-0. xxvii + 607 pp. LCCN QA76.76.O63 L487 1991b. US$34.95. March 1994 printing with corrections, updates, and December 1991 Appendix G.

NISO:1991:EMP

[Org91] National Information Standards Organization. Electronic Manuscript Preparation and Markup: American National Standard for Electronic Manuscript Preparation and Markup. Transaction Publishers, Rutgers University, New Brunswick, NJ 08903, USA, 1991. ISBN 0-88738-945-7. xv + 167 pp. LCCN Z286.E43 E428 1991. ANSI/NISO Z39.59-1988. Approved December 1, 1988, by American National Standards Institute. Developed by The National Information Standards Organization.

IEEE:p754

[P7585] IEEE Task P754. ANSI/IEEE 754-1985, Standard for Binary Floating-Point Arithmetic. IEEE, New York, August 12 1985. A preliminary draft was published in the January 1980 issue of IEEE Computer, together with several companion articles. Available from the IEEE Service Center, Piscataway, NJ, USA.

Plauger:SCL92

[Pla92] P. J. Plauger. The Standard C Library. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-838012-0. xiv + 498 pp. LCCN QA76.73.C15 P563 1991.

Plum:NDC87

[Plu87] T. Plum. Notes on the Draft C Standard. Plum Hall, 1 Spruce Ave, Cardiff, NJ 08232, USA, 1987. ISBN 0-911537-06-6. 92 pp. LCCN QA76.73.C15 P585 1987. US$10.00.

IEEE:POSIX.1-90

[POS90] System Application Program Interface (API) [C Language]. Information technology—Portable Operating System Interface (POSIX). IEEE Computer Society, 345 E. 47th St, New York, NY 10017, USA, 1990. ISBN 1-55937-061-0. LCCN 90-084554.

IEEE:POSIX.2-93-vol1

[POS93a] Portable Operating System Interface (POSIX)—Part 2: Shell and Utilities (Volume 1). Information technology—Portable Operating System Interface (POSIX). IEEE Computer Society, 345 E. 47th St, New York, NY 10017, USA, 1993. ISBN 1-55937-255-9. 1–750 pp.

IEEE:POSIX.2-93-vol2

[POS93b] Portable Operating System Interface (POSIX)—Part 2: Shell and Utilities (Volume 2). Information technology—Portable Operating System Interface (POSIX). IEEE Computer Society, 345 E. 47th St, New York, NY 10017, USA, 1993. ISBN 1-55937-255-9. 751–1195 pp.

Schildt:1990:AAC

[Sch90] Herbert Schildt. The Annotated ANSI C Standard: American National Standard for Programming Languages: C. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881952-0. xvi + 219 pp. LCCN QA76.73.C15S356 1990. US$39.95.

SIGGRAPH:core77

[SIG77] ACM/SIGGRAPH. Status report of the Graphic Standards Planning Committee of ACM/ SIGGRAPH. ACM SIGGRAPH— Computer Graphics, 11(3), 1977.

SIGGRAPH:core79

[SIG79] ACM/SIGGRAPH. Status report of the Graphic Standards Planning Committee of ACM/ SIGGRAPH. ACM SIGGRAPH— Computer Graphics, 13(3), August 1979.

Smith:SGM86

[Smi87a] Joan M. Smith. The Standard Generalized Markup Language and Related Issues. British National Bibliography Research Fund, London, UK, 1987. ISBN 0-7123-3082-8. 63 pp. LCCN Z286 .E43 S65 1986.

Smith:SGM87

[Smi87b] Joan M. Smith. The Standard Generalized Markup Language (SGML): guidelines for authors. British Library Publications Sales Unit, Wolfeboro, NH, USA, 1987. ISBN 0-7123-3112-3. ix + 66 pp. LCCN QA76.73.S44 S65x 1987.

Smith:SRS92

[Smi92] Joan M. Smith. SGML and Related Standards: Document Description and Processing Languages. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-806506-3. xviii + 151 pp. LCCN QA76.73 .S42 1992. US$50.25.

Smith:1988:SUG

[SS88] Joan M. Smith and Robert Stutely.

SGML: The User’s Guide to ISO 8879. Ellis Horwood series in computers and their applications. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, 1988. ISBN 0-470-21126-1 (Halsted Press), 0-7458-0221-4. 175 pp. LCCN QA76.73.S44 S441 1988. UK£22.50.

Stevenson:COMPUTER-14-3-51

[Ste81] David Stevenson. A proposed standard for binary floating-point arithmetic. Computer, 14(3):51– 62, March 1981.

Summers:1993:IIW

[Sum93] Clayton Summers. Introduction to ISO 9660: what it is, how it is implemented, and how it has been extended. Disc Manufacturing, 1409 Foulk Rd., Suite 102, Wilmington, DE 19803, USA, 1993. 29 + vi pp.

Unicode:1991:USW

[The91] The Unicode Consortium. The Unicode Standard: Worldwide Character Encoding. Version 1.0. Volumes 1 and 2. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-56788-1 (paperback, vol. 1), 0-201-60845-6 (paperback, vol. 2). LCCN QA268 .U55 1991, Z103 .U6 1991. US$39.95 (paperback), US$32.95 (hardcover).

ANSI:ftn69

[X3J69] ANSI Subcommittee X3J3. Clarification of Fortran standards— initial progress. Communications of the Association for Computing Machinery, 12:289–294, 1969. See also [Ame66].

ANSI:ftn71

[X3J71] ANSI Subcommittee X3J3. Clarification of Fortran standards— second report. Communications of the Association for Computing Machinery, 14:628–642, 1971. See also [Ame66].

ANSI:ftn76

[X3J76] ANSI Subcommittee X3J3. Draft proposed ANS Fortran. ACM SIG-PLAN Notices, 11(3), 1976. See also final standard [Ame78].

Zlotnick:PSP-91

[Zlo91] Fred Zlotnick. The POSIX.1 Standard: A Programmer’s Guide. Benjamin/Cummings, 390 Bridge Pkwy., Redwood City, CA 94065, USA, 1991. ISBN 0-8053-9605-5. LCCN QA76.76.O63 Z57.